OpenStudy (cieloveswork):

Complete the solution table from left to right for the quadratic function y=x^2-11x+30 x|3|4|5|6| y|? |? |? |?| A. 0, 0, 14, 37 B. −9, 16, 25, 0 C. −4, 5, −6, 0 D. 6, 2, 0, 0

TheSmartOne (thesmartone):

plug in x = 3 into the quadratic function and solve for y plug in x = 4 into the quadratic function and solve for y plug in x = 5 into the quadratic function and solve for y plug in x = 6 into the quadratic function and solve for y tell me what you get

OpenStudy (cieloveswork):

b....idk

TheSmartOne (thesmartone):

plug in x = 3 y=x^2-11x+30 y = (3)^2 - (11 * 3) + 30 y = ??

OpenStudy (cieloveswork):

y=6

TheSmartOne (thesmartone):

which answer choice starts off with 6?

OpenStudy (cieloveswork):

d

TheSmartOne (thesmartone):

we don't even have to do the rest of the plugging in since only one answer choice starts with it correct :)

OpenStudy (cieloveswork):

thx u

TheSmartOne (thesmartone):

otherwise we would have to plug in x = 4, and x = 5, and x = 6 basically you replace x with the number and solve :)

TheSmartOne (thesmartone):

Anytime! :D

OpenStudy (skullpatrol):

Correct, notice that only one answer choice has a value of 6 ;-)

OpenStudy (skullpatrol):

Oops sorry. This site lags ALOT :-/

OpenStudy (skullpatrol):

Even though it is NOW "Powered by Brainly" :)

OpenStudy (skullpatrol):

In fact, the value of 6 is defined as "the value of the function" since it is a member of the range of the function.